# Trajectory Forecasting through Low-Rank Adaptation of Discrete Latent Codes

### v1

@@ 1 Introduction
Pedestrian trajectory forecasting finds applications in video surveillance [19], behavioural analysis [28], intrusion detection [33], and autonomous systems [3]. The goal is to predict the future paths of a set of agents from a few observations of their motion. The prediction can incorporate the interactions between pedestrians [15, 32, 23], or visual attributes of the environment they move within [6].
As multiple plausible paths can be forecast, trajectory prediction reveals an uncertain and multi-modal nature. To achieve this, recent data-driven approaches [22, 12, 16] lean toward a stochastic formulation that places a distribution  over  the  future  trajectory,  rather  than  a  single  estimated  path  with  $100\%$
certainty (i.e.deterministic approaches [23]). In doing so, recent stochastic methods take advantage of the latest breakthroughs in deep generative modeling for image generation. For example, [12, 29] resorted to Generative Adversarial Networks, while [44, 16, 35] borrowed ideas from the class of variational methods.
@@
In the field of image generation, Vector Quantized Variational Autoencoders [39] (VQ-VAEs) have demonstrated to mitigate posterior collapse. VQ-VAEs models avoid the hand-crafted Gaussian prior distribution; differently, they build upon a learnable categorical prior, thereby yielding a discrete latent space. The symbols of this space are the keys of a fixed-size dictionary (codebook), whose values are learnable latent codes. Thanks to the resulting increased flexibility, VQ-VAEs embody a promising (yet still unexplored) paradigm for trajectory forecasting.
@@ 4.2 Instance-based Codebook
Our idea is to modify the content of the codebook, such that it reflects the motion observed in the past trajectory. The intuition is that different motion styles (e.g., straight vs. curvilinear) could prefer distinct latent codes and discretization strategies. On this basis, we exploit again the contextual features $h_{\operatorname{ctx}}$ to generate an instance-based codebook $\xi=f_{\xi}(\cdot,h_{\operatorname{ctx}})$, computed through a tailored learnable module $f_{\xi}$. The latter shares the same design of the above-described encoding networks and hence builds upon social-temporal transformers [23]. Afterward, we combine static and instance-based codebooks by means of summation, thus obtaining a conditioned codebook $e_{c}$:
@@
Moreover, the way we define the codebook draws inspiration from the successes of low-rank adaptation [14] for finetuning Large Language Models (LLMs). Namely, we opt for a low-rank characterization of $f_{\xi}$, which means that the instance-driven modifications to the static codebook lie on a lower-dimensional manifold of the parameter space. We hence define the instance-based codebook $\xi$ as a matrix product of two low-rank matrices $B_{\operatorname{ctx}}$ and $A$, as follow:
@@
$\xi$ $=B_{\operatorname{ctx}}A\quad\text{where }A\in\mathbb{R}^{r\times k}.$
@@ 5 Experiments
Implementation Details. We set the number of codewords $C$ to $16$ for all datasets, while we take the best rank $r$ for each dataset (e.g., 8 for SDD and NBA, 4 for NFL). For the first stage, we use AdamW [21] as optimizer with $\operatorname{lr}=5\times 10^{-4}$, $\beta_{1}=0.5$ and $\beta_{2}=0.9$. We train on SDD for $7000$ epochs with batch size equal to $256$. For NBA and NFL, we instead optimize for $700$ epochs (the batch size equals $64$). We use a cosine schedule for $\lambda_{d}$ from an initial value of 00 to a final value of $1$. In this way, we can introduce the instance-level codebook gradually during training.
@@
We will make the code publicly available upon acceptance.
@@ 5.1 On the Impact of the Instance-based Codebook
Tab. 2 presents the evaluation results for different ranks $r$. We observe that a higher reconstruction capability during the initial training stage is associated with increased difficulty in the diffusion task, resulting in lower accuracy. This indicates a correlation between the two phases: achieving optimal results in the first phase does not necessarily yield the best final generation metrics, as it complicates the joint task of trajectory generation (i.e., sampling from the prior and reconstructing through the decoder). Tab. 2 demonstrates that the most favorable final metrics are achieved by striking a balance between low reconstruction error and good diffusion accuracy. All metrics reported in the two tables are computed on the test set.
@@ 5.2 Comparison with SOTA Methods
On the SDD dataset (see Tab. 3), we attain superior results in terms of ADE and closely match the FDE achieved by MemoNet. While PECNet and GroupNet, among C-VAE methods, demonstrate noteworthy performance compared to the older S-GAN and Trajectron++, they struggle in terms of FDE, especially when compared to MemoNet. This could be ascribed to the effective sampling strategy of MemoNet, which integrates a tailored clustering phase to generate multiple overall intentions.
@@ 5.3 Qualitative Results
6 Conclusion
We introduce a novel stochastic approach for trajectory prediction. It builds upon Vector Quantization to yield a predictive distribution that preserves both sampling fidelity and diversity. Our main contribution lies in the creation of a dynamic, instance-related codebook, allowing it to reflect past trajectory information. Notably, contextual information is incorporated into the codebook through a low-rank update. We conduct several empirical studies to validate our approach, demonstrating its superior generative capabilities compared to both standard VQ-VAEs and existing methods. This leads to state-of-the-art results on three established benchmarks.

### v2

@@ 1 Introduction
Trajectory forecasting finds applications in video surveillance [19], multi-object tracking [22, 6], behavioural analysis [29], and intrusion detection [34]. The goal is to predict the future paths of a set of agents from a few observations of their motion. The prediction can incorporate the interactions between pedestrians [15, 33, 24], or visual attributes of the environment they move within [5].
As multiple plausible paths can be forecast, trajectory prediction reveals an uncertain and multi-modal nature. To achieve this, recent data-driven approaches [23, 12, 16] lean toward a stochastic formulation that places a distribution over the future trajectory, rather than a single estimated path with $100\%$ certainty (i.e.deterministic approaches [24]). In doing so, recent stochastic methods take advantage of the latest breakthroughs in deep generative modeling for image generation. For example, [12, 30] resorted to Generative Adversarial Networks, while [45, 16, 36] borrowed ideas from the class of variational methods.
@@
In the field of image generation, Vector Quantized Variational Autoencoders [40] (VQ-VAEs) have proven to mitigate posterior collapse. VQ-VAEs models avoid the hand-crafted Gaussian prior distribution; differently, they build upon a learnable categorical prior, thereby yielding a discrete latent space. The symbols of this space are the keys of a fixed-size dictionary (codebook), whose values are learnable latent codes. Thanks to the resulting increased flexibility, VQ-VAEs embody a promising paradigm for trajectory forecasting.
@@ 4.2 Instance-based Codebook
Our idea is to modify the content of the codebook, such that it reflects the motion observed in the past trajectory. The intuition is that different motion styles (e.g., straight vs. curvilinear) could prefer distinct latent codes and discretization strategies. On this basis, we exploit again the contextual features $h_{\operatorname{ctx}}$ to generate an instance-based codebook $\xi=f_{\xi}(\cdot,h_{\operatorname{ctx}})$, computed through a tailored learnable module $f_{\xi}$. The latter shares the same design of the above-described encoding networks and hence builds upon social-temporal transformers [24]. Afterwards, we combine static and instance-based codebooks by means of summation, thus obtaining a conditioned codebook $e_{c}$:
@@
Moreover, the way we define the codebook draws inspiration from the successes of low-rank adaptation [14] for fine-tuning Large Language Models (LLMs). Namely, we opt for a low-rank characterization of $f_{\xi}$, which means that the instance-driven modifications to the static codebook lie on a lower-dimensional manifold of the parameter space. We hence define the instance-based codebook $\xi$ as a matrix product of two low-rank matrices $B_{\operatorname{ctx}}$ and $A$, as follow:
@@
$\xi$ $=B_{\operatorname{ctx}}A\quad\text{where }A\in\mathbb{R}^{r\times C}.$
@@ 5 Experiments
Implementation Details [^1] The code is available at https://github.com/aimagelab/LRVQ.. We set the number of codewords $C$ to $16$ for all datasets, while we take the best rank $r$ for each dataset (e.g., 8 for SDD and NBA, 4 for NFL). For the first stage, we use AdamW [21] as optimizer with $\operatorname{lr}=5\times 10^{-4}$, $\beta_{1}=0.5$ and $\beta_{2}=0.9$. We train on SDD for $7000$ epochs with batch size equal to $256$. For NBA and NFL, we instead optimize for $700$ epochs (the batch size equals $64$). We use a cosine schedule for $\lambda_{\xi}$ from an initial value of 00 to a final value of $1$. In this way, we can introduce the instance-level codebook gradually during training.
@@ 5.1 On the Impact of the Instance-based Codebook
Tab. 2 presents the results for different ranks $r$. We observe that a higher reconstruction capability during the initial training stage is associated with increased difficulty in the diffusion task, resulting in lower accuracy. This indicates a correlation between the two phases: achieving optimal results in the first phase does not necessarily yield the best final generation metrics, as it complicates the joint task of trajectory generation (i.e., sampling from the prior and reconstructing through the decoder). Tab. 2 demonstrates that the most favorable final metrics are achieved by striking a balance between low reconstruction error and good diffusion accuracy.
@@ 5.2 Comparison with SOTA Methods
On the SDD dataset (Tab. 3), we attain superior ADE results, matching closely MemoNet in FDE. While PECNet and GroupNet, among C-VAE methods, demonstrate noteworthy performance compared to the older S-GAN and Trajectron++, they struggle in FDE, especially when compared to MemoNet. This could be ascribed to the effective sampling strategy of MemoNet, which integrates a tailored clustering phase to generate multiple overall intentions.
@@ 5.3 Qualitative Results
6 Discussion and Conclusions
Limitations. The complexity of our model is linked to two factors:
  • Two-step training procedure: although VQ-VAE offers benefits such as a learned prior, the training must be divided into two distinct stages, which increases the total time required to train the model.
  • Inference time: the inference procedure described in Sec. 4.4 takes longer as the number $N$ of starting guesses increases. To obtain a trade-off between the accuracy of the ensemble of $K$ final generations and the computational time, the parameter $N$ has to be carefully adjusted.
Conclusion. We propose a stochastic approach for trajectory prediction. It builds upon Vector Quantization to yield a predictive distribution that preserves both sampling fidelity and diversity. Our main contribution lies in a dynamic, instance-related codebook encompassing past trajectory information. Notably, contextual information is incorporated into the codebook through a low-rank update. We conduct several empirical studies to validate our approach, demonstrating its superior generative capabilities compared to both standard VQ-VAEs and existing methods. This leads to state-of-the-art results on three established benchmarks.
Acknowledgement
The research was supported by the Italian Ministry for University and Research through the PNRR project ECOSISTER ECS 00000033 CUP E93C22001100001 and by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237).